# SketchOGD: Memory-Efficient Continual Learning

### v2

@@ I Introduction
Humans excel at learning continuously, acquiring new capabilities throughout their lifespan without repeatedly relearning previously mastered tasks. Moreover, knowledge gained in one context can often transfer to other tasks, and human learning remains relatively robust in the face of natural distribution shifts [1, 2]. In contrast, endowing deep neural networks with similar continual learning abilities remains a challenge, largely due to catastrophic forgetting [3]. Under the naïve approach of training networks on tasks sequentially, performance on earlier tasks can decline precipitously once new tasks are introduced.
@@
To advance memory-efficient continual learning, we center our attention on a conceptually straightforward yet powerful continual learning algorithm known as orthogonal gradient descent (OGD) [4]. Building on the properties of overparameterized neural networks [5, 6, 7, 8], OGD preserves performance on previous tasks by storing model gradients for previous training examples and projecting later weight updates into the subspace orthogonal to those gradients. We choose to focus on OGD because it is simple, effective, and amenable to analysis. Furthermore, unlike other continual learning algorithms such as A-GEM [9], it does not require re-computing forward passes on old tasks, meaning that there is no need to store raw datapoints. This reduces computation requirements and may be beneficial in a setting where privacy is important. However, a key limitation of OGD is the memory cost of storing large numbers of gradients. For a model with $p$ parameters trained for $T$ iterations, the memory requirement is $\mathcal{O}(pT)$. For deep overparameterized networks, $p$ itself may be large, and the growth in the memory requirement with $T$ makes OGD ill-suited to learning over long time horizons.
To resolve the memory issues of OGD, this paper proposes a memory-efficient continual learning algorithm called sketched orthogonal gradient descent—or SketchOGD. More specifically, we leverage sketching algorithms [10] to compress the past model gradients required by OGD into a lower-dimensional representation known as a sketch. Crucially, we can update the sketch efficiently (without needing to decompress and recompress gradients already encountered) such that SketchOGD can run online. In short, SketchOGD avoids storing the full history of past gradients by maintaining a sketch of a fixed size that may be updated cheaply and online. To ensure that SketchOGD is a viable continual learning algorithm, we need to characterize how effective the sketching is in compressing past gradients. We tackle this question both theoretically and experimentally. More specifically, our contributions are as follows:
@@ I-A Related Work
Research on lifelong and continual learning focuses on solving the problem of catastrophic forgetting [3, 11]. Some approaches are based on rehearsing old data/experiences whilst learning new tasks [12, 13, 14, 15, 16, 17] or allocating neural resources incrementally for new tasks [18, 19, 20, 21], while others rely on regularizing training to preserve performance on earlier tasks [22]. Orthogonal gradient descent is one such regularization-based approach, which we discuss in Section II-A. Other methods include elastic weight consolidation (EWC) [22], synaptic intelligence [23] and averaged gradient episodic memory (A-GEM) [9]. Parisi et al. [24] provide a more thorough review of recent developments. A variant of OGD, known as ORFit, was recently proposed by Min et al. [25] for the special setting of one-pass learning.
@@
Many approaches have been proposed for approximating or compressing a matrix, including techniques based on sparsification and quantization [26], column selection [27], dimensionality reduction [28] and approximation by submatrices [29]. One practical form of dimensionality reduction known as sketching is introduced in Section II-B. Halko et al. [28] provide a brief survey of the matrix approximation literature. Sketching has also been successfully employed for summarizing the curvature of neural networks for the purpose of out-of-distribution detection [30], motivating its application to OGD.
@@ II-A Orthogonal Gradient Descent
To be more precise, given a model $f$ with weights $w\,{\in}\,\mathbb{R}^{p}$, OGD stores the gradients $\nabla_{w}f(x_{1},w),...,\nabla_{w}f(x_{n},w)$ for training examples $x_{1},...,x_{n}$. This paper focuses on classification, and only the gradients of the ground-truth logit are stored which has been found to improve performance empirically [4]. The $n$ gradient vectors are arranged into a matrix $G\in\mathbb{R}^{p\times n}$, and future weight updates are projected onto the orthogonal subspace $(\operatorname{range}G)^{\perp}$. To perform the projection, one may form an orthonormal basis for $\operatorname{range}G$. Letting $\operatorname{orth}G$ denote a matrix whose columns form such a basis, then the OGD weight update may be written as
@@
A major limitation of OGD is the memory requirement of storing all $n$ past gradients, where the number of trained examples $n$ grows with the runtime of the algorithm. A simple solution is to only store a random subset of past gradients, in the hope that the corresponding submatrix of $G$ will accurately approximate the range of the full $G$ matrix. We call this method RandomOGD.
Another solution called PCA-OGD has been proposed [31], which stores only the top principal components of the gradients for each task. Similarly, ORFit [25] uses incremental PCA to derive a variant of OGD that is specialized for one-pass learning. Incremental PCA could be directly applied to OGD to compress the memory to a fixed size in an online manner. However, its choice of which principal components to discard, and thus, the quality of compression, is sensitive to the order in which the training data is encountered. The next section introduces an alternative means of approximation that avoids this issue.
II-B Matrix Approximation via Sketching
To reduce the memory cost of OGD, we use sketch, which approximates a matrix in lower dimensions:
@@ II-B Matrix Approximation via Sketching
Given a sketch $(\Omega,\Psi,Y,W)$ of a matrix $A\in\mathbb{R}^{m\times n}$, and letting $Q=\operatorname{orth}Y\in\mathbb{R}^{m\times k}$, $\widehat{A}\vcentcolon=Q(\Psi Q)^{\dagger}W$ approximates $A$.
Rigorous justification for this is given by Tropp et al. [10]. The quality of the approximation improves with the size of the sketch set by $k$ and $l$. Especially, $k$ determines the rank of $\widehat{A}$.
@@
Given a sketch $(\Omega,\Psi,Y,W)$ of a symmetric matrix $A\in\mathbb{R}^{m\times m}$, let $Q=\operatorname{orth}Y\in\mathbb{R}^{m\times k}$ and define $X\vcentcolon=(\Psi Q)^{\dagger}W\in\mathbb{R}^{k\times m}$. Then, $\widehat{A}_{\mathrm{sym}}\vcentcolon=\tfrac{1}{2}(QX+X^{\top}Q^{\top})$ approximates $A$.
Formal justification is again given by Tropp et al. [10]. To approximate $\operatorname{range}A$, we can use the range of the concatenation $[Q\;X^{\top}]$ from this proposition as we observe that:
@@
III Sketched Orthogonal Gradient Descent
This section applies the matrix sketching methods introduced in Section II-B to obtain three variants of sketched orthogonal gradient descent (SketchOGD). We propose different variants as it is a priori unclear which performs best, but we will investigate this both by presenting theoretical bounds (Section IV-B) and by running experiments (Section V). All variants adopt the following pattern:
  1. 1.
  2. 2.
  3. 3.
Algorithm 1 SketchOGD. The algorithm requires a model with weight vector $w$, a base optimizer such as SGD, as well as sketching subroutines from Sketching Method 2, 3 or 4.
@@ III Sketched Orthogonal Gradient Descent
Sketching Method 2 Sketch of gradient matrix $G$.
@@
Sketching Method 3 Sketch of product $GG^{\top}$.
@@
Sketching Method 4 Symmetric sketch of product $GG^{\top}$.
@@
This procedure is described formally in Algorithm 1. Three variants are obtained by using either Sketching Method 2, 3 or 4 to supply the subroutines $\mathtt{InitializeSketch}$, $\mathtt{UpdateSketch}$ and $\mathtt{ExtractBasisFromSketch}$. The variants arise since there is freedom in both which matrix is sketched and what sketching technique is used. For instance, since $G$ and $GG^{\top}$ have the same range, either matrix may be sketched to approximate $\operatorname{range}G$. Furthermore, since $GG^{\top}$ is symmetric, Proposition II.3 may be applied. To elaborate, we propose three variants of SketchOGD:
SketchOGD-1: Algorithm 1 + Sketching Method 2. This algorithm directly sketches the gradient matrix $G$ via Proposition II.2, and $\operatorname{orth}Y$ for $Y=G\Omega$ forms the orthonormal basis. When a new gradient $g$ is received, we update $Y$ online instead of recomputing it from scratch since $Y^{(new)}=G^{(new)}\Omega^{(new)}=[G\;g]\begin{bmatrix}\Omega\\ \omega^{\top}\end{bmatrix}=Y+g\omega^{\top}$. Here, the random matrix $\Omega$ is incremented with a standard normal vector $\omega$ since the dimension of $G$ is changed. In terms of memory, SketchOGD-1 needs to store only $Y$ at a cost of $pk$.
SketchOGD-2: Algorithm 1 + Sketching Method 3. This algorithm sketches the product $GG^{\top}$ via Proposition II.2 as in SketchOGD-1. To update the sketch variable $Y=GG^{\top}\Omega$ online when a new gradient $g$ is received, we leverage the decomposition $Y^{(new)}=G^{(new)}{G^{(new)}}^{\top}\Omega=\sum_{c\in G}cc^{\top}\Omega=Y+gg^{\top}\Omega$ where the summation is running over the columns of $G$. In terms of memory, SketchOGD-2 needs to store both $Y$ and $\Omega$ at a cost of $2pk$.
SketchOGD-3: Algorithm 1 + Sketching Method 4. This algorithm sketches the product $GG^{\top}$ via symmetric sketch (Proposition II.3). The sketch variables $Y$ and $W$ are updated online using the decomposition of $GG^{\top}$ as in SketchOGD-2. The computation of the pseudoinverse $(\Psi Q)^{\dagger}$ is expedited by first taking the orthogonal-triangular ($\mathtt{QR}$) decomposition. The full sketch needs to be stored, at a memory cost of $2p(l+k)$. Since $l\geq k$, this is at least four times that of SketchOGD-1. However, the basis returned by Sketching Method 4 spans a higher-dimensional subspace.
@@ IV-A Constructing a Metric Suited to SketchOGD
Given a matrix $G$ to be approximated and a matrix $B$ returned by a sketching method, the reconstruction error is given by:
$\mathcal{E}_{G}(B)=\sum_{g\in G}\|g-\operatorname{proj}_{B}g\|_{2}^{2}=\|(1-\operatorname{proj}_{B})G\|_{F}^{2},$
where the summation is running over the columns of G. In SketchOGD, if the metric vanishes while $\operatorname{range}B\subseteq\operatorname{range}G$, then SketchOGD is equivalent to OGD.
@@
Another possible metric is the Grassmann distance between subspaces [32], which could be applied to measure the distance between the subspace $\operatorname{range}G$ and the subspace $\operatorname{range}B$, where $B$ is returned by the sketching method. The drawback to this approach is that it neglects how the columns $\{g\in G\}$ are distributed within $\operatorname{range}G$. If certain directions appear more frequently, then it is more important for OGD that these directions are well-modeled.
@@ IV-B Deriving Bounds
We now provide a novel theoretical analysis of sketching applied to continual learning, through bounding our metric. While Tropp et al. [10] analyzed the Frobenius reconstruction error of sketching $\mathbb{E}\|A-\widehat{A}\|_{F}^{2}$, and Doan et al. [31] bounded catastrophic forgetting, to our knowledge we are the first to theoretically analyze the disruption a compression method causes to OGD, specifically.
@@
Define $U\in\mathbb{R}^{p\times p},\Sigma_{G}\in\mathbb{R}^{p\times n},V\in\mathbb{R}^{n\times p}$ as the singular value decomposition (SVD) of $G$ such that $G=U\Sigma_{G}V$, where $n$ is the number of sketched gradients. Then, we can define $\Sigma\vcentcolon=\Sigma_{G}\Sigma_{G}^{\top}\in\mathbb{R}^{p\times p}$ such that $GG^{\top}=U\Sigma U^{\top}$. Assuming singular values sorted in decreasing magnitude, we define $\Sigma_{1},\Sigma_{2},U_{1},U_{2}$ by splitting matrices at an index $\gamma\leq k-2$:
@@
Theoretical bounds on our metric based on this split would imply their dependence on the singular value decay of the gradients. In the basic sketching method where we directly sketch the matrix of gradients $G$, we can adapt a theorem of [28] as below. See the appendix for the proof.
Theorem IV.1 (Expected error in Sketching Method 2).
Given that we create a sketch $\widehat{G}$ of $G$ with rank $k$ in accordance with Sketching Method 2, and use the setup described in Definition IV.2, then the expected value of the metric $\mathcal{E}_{G}(\widehat{G})$ has an upper bound: $\mathbb{E}[\mathcal{E}_{G}(\widehat{G})]\leq\min_{\gamma\leq k-2}(1+\tfrac{\gamma}{k-\gamma-1})\cdot\operatorname{Tr}{\Sigma_{2}}.$
@@
Note that sharper decays push the optimal split location $\gamma$ away from 00. For intuition, consider a flat spectral decay such that $\Sigma=\lambda I_{p\times p}$ for some $\lambda>0$ and the identity matrix $I_{p\times p}$. Then, $\gamma=0$ attains the minimum of the upper bound $p\lambda$. On the other hand, if $\Sigma=\Sigma_{G}\Sigma_{G}^{T}$ has linearly decaying spectral decay from $2\lambda$ to 00, then $\gamma=2(k{-}1){-}p$ attains the minimum approximately at $\frac{4\lambda}{p}(k{-}1)(p{-}k{+}1)$ for large enough $k$. Analysis can be found in the appendix.
We now provide two novel theorems for Sketching Method 3, one upper bounding the metric deterministically and the other in expectation. See the appendix for the proofs.
Theorem IV.2 (Deterministic error in Sketching Method 3).
We define $G,\Sigma_{1},\Sigma_{2},U_{1},U_{2}$ as in Definition IV.2 for a sketch parameter $k$ and split position $\gamma\leq k-2$. If we create a sketch $(\Omega,\Psi,Y,W)$ and reconstruction $\widehat{A}$ of $A=GG^{\top}$ in accordance with Sketching Method 3, we can use $\Omega\in\mathbb{R}^{p\times k}$ to define: $\Omega_{1}:=U_{1}^{\top}\Omega$, $\Omega_{2}:=U_{2}^{\top}\Omega$. Assume $\Omega_{1}\in\mathbb{R}^{\gamma\times k}$ is full-rank. Then, if $\gamma$ is greater than the rank of $G$, then the metric $\mathcal{E}_{G}(\widehat{A})=0$. Else, $\mathcal{E}_{G}(\widehat{A})\leq\|\Sigma_{2}\Omega_{2}\Omega_{1}^{\dagger}\Sigma_{1}^{-1/2}\|_{F}^{2}+\|\Sigma_{2}^{1/2}\|_{F}^{2}$.
@@
Theorem IV.3 (Exptected error in Sketching Method 3).
We define $G,\Sigma_{1},\Sigma_{2},U_{1}$ as in Definition IV.2 for a sketch parameter $k$ and arbitrary split position $\gamma\leq k-2$. We create a sketch $(\Omega,\Psi,Y,W)$ and reconstruction $\widehat{A}$ of $A=GG^{\top}$, and define $\Omega_{1}:=U_{1}^{\top}\Omega$ as in Theorem IV.2. We assume $\Omega_{1}$ has full rank. Then, if $\gamma$ is greater than the rank of $G$, then the metric $\mathbb{E}[\mathcal{E}_{G}(\widehat{A})]=0$. Else,
$\mathbb{E}[\mathcal{E}_{G}(\widehat{A})]\leq\min_{\gamma\leq k-2}\tfrac{\gamma}{k-\gamma-1}\operatorname{Tr}(\Sigma_{2}^{2})\operatorname{Tr}(\Sigma_{1}^{-1})+\operatorname{Tr}(\Sigma_{2}).$

### v3

@@ I Introduction
Humans excel at learning continuously, acquiring new capabilities throughout their lifespan without repeatedly relearning previously mastered tasks. Moreover, knowledge gained in one context can often transfer to other tasks, and human learning remains relatively robust in the face of natural distribution shifts [barnett2002and, bremner2012multisensory]. In contrast, endowing deep neural networks with similar continual learning abilities remains a challenge, largely due to catastrophic forgetting [mccloskey]. Under the naïve approach of training networks on tasks sequentially, performance on earlier tasks can decline precipitously once new tasks are introduced.
@@
To advance memory-efficient continual learning, we center our attention on a conceptually straightforward yet powerful continual learning algorithm known as orthogonal gradient descent (OGD) [Farajtabar2019OrthogonalGD]. Building on the properties of overparameterized neural networks [li2018learning, azizan2018stochastic, azizan2019stochastic, allen2019convergence], OGD preserves performance on previous tasks by storing model gradients for previous training examples and projecting later weight updates into the subspace orthogonal to those gradients. We choose to focus on OGD because it is simple, effective, and amenable to analysis. Furthermore, unlike other continual learning algorithms such as A-GEM [Chaudhry2018EfficientLL], it does not require re-computing forward passes on old tasks, meaning that there is no need to store raw datapoints. This reduces computation requirements and may be beneficial in a setting where privacy is important. However, a key limitation of OGD is the memory cost of storing large numbers of gradients. For a model with $p$ parameters trained for $T$ iterations, the memory requirement is $\mathcal{O}(pT)$. For deep overparameterized networks, $p$ itself may be large, and the growth in the memory requirement with $T$ makes OGD ill-suited to learning over long time horizons.
To resolve the memory issues of OGD, this paper proposes a memory-efficient continual learning algorithm called sketched orthogonal gradient descent—or SketchOGD. More specifically, we leverage sketching algorithms [Tropp] to compress the past model gradients required by OGD into a lower-dimensional representation known as a sketch. Crucially, we can update the sketch efficiently (without needing to decompress and recompress gradients already encountered) such that SketchOGD can run online. In short, SketchOGD avoids storing the full history of past gradients by maintaining a sketch of a fixed size that may be updated cheaply and online. To ensure that SketchOGD is a viable continual learning algorithm, we need to characterize how effective the sketching is in compressing past gradients. We tackle this question both theoretically and experimentally. More specifically, our contributions are as follows:
@@ I-A Related Work
Research on lifelong and continual learning focuses on solving the problem of catastrophic forgetting [mccloskey, Ratcliff1990ConnectionistMO]. Some approaches are based on rehearsing old data/experiences whilst learning new tasks [rehearsal, shin2017continual, rebuffi2017icarl, rolnick2019experience, van2020brain, pellegrini2020latent] or allocating neural resources incrementally for new tasks [Rusu2016ProgressiveNN, valkov2018houdini, veniat2021efficient, ostapenko2021continual], while others rely on regularizing training to preserve performance on earlier tasks [Kirkpatrick2016OvercomingCF]. Orthogonal gradient descent is one such regularization-based approach, which we discuss in Section II-A. Other methods include elastic weight consolidation (EWC) [Kirkpatrick2016OvercomingCF], synaptic intelligence [Zenke2017ContinualLT] and averaged gradient episodic memory (A-GEM) [Chaudhry2018EfficientLL]. PARISI201954 provide a more thorough review of recent developments. A variant of OGD, known as ORFit, was recently proposed by ORFit for the special setting of one-pass learning.
@@
Many approaches have been proposed for approximating or compressing a matrix, including techniques based on sparsification and quantization [achlioptas], column selection [selection], dimensionality reduction [halko] and approximation by submatrices [submatrix]. One practical form of dimensionality reduction known as sketching is introduced in Section II-C. halko provide a brief survey of the matrix approximation literature. Sketching has also been successfully employed for summarizing the curvature of neural networks for the purpose of out-of-distribution detection [sharma2021sketching], motivating its application to OGD.
@@ II-A Orthogonal Gradient Descent
To be more precise, given a model $f$ with weights $w\,{\in}\,\mathbb{R}^{p}$, OGD stores the gradients $\nabla_{w}f(x_{1},w),...,\nabla_{w}f(x_{n},w)$ for training examples $x_{1},...,x_{n}$. This paper focuses on classification, and only the gradients of the ground-truth logit are stored which has been found to improve performance empirically [Farajtabar2019OrthogonalGD]. The $n$ gradient vectors are arranged into a matrix $G\in\mathbb{R}^{p\times n}$, and future weight updates are projected onto the orthogonal subspace $(\operatorname{range}G)^{\perp}$. To perform the projection, one may form an orthonormal basis for $\operatorname{range}G$. Letting $\operatorname{orth}G$ denote a matrix whose columns form such a basis, then the OGD weight update may be written as
@@
II-B Memory Compression for OGD
A major limitation of OGD is the memory requirement of storing all $n$ past gradients, where the number of trained examples $n$ grows with the runtime of the algorithm. A simple solution is to only store a random subset of past gradients, in the hope that the corresponding submatrix of $G$ will accurately approximate the range of the full $G$ matrix. We call this method RandomOGD. Another solution called PCA-OGD has been proposed [Doan2020ATA], which stores only the top principal components of the gradients for each task. Similarly, ORFit [ORFit] uses incremental PCA to derive a variant of OGD that is specialized for one-pass learning.
The core observation behind compressing the gradient matrix $G$ is that not all directions in $G$ are equally important. Figures 2 and 3 show that the singular values of the gradient matrices exhibit strong decay, so most of the “energy” of $G$ is concentrated in a relatively low-dimensional subspace. This suggests that many directions in $\operatorname{range}G$ rarely contribute to preventing catastrophic forgetting across tasks. By focusing on the dominant directions, we can retain the part of the subspace that contributes most to OGD’s core mechanism.
However, classical PCA used in PCA-OGD requires storing all gradients for a task to perform compression, which introduces a memory overhead that grows with task size and requires knowing task boundaries. Also, its memory to store the compressed results grows with each task, which makes it hard to predict memory costs without knowing the number of tasks in advance.
Incremental PCA [artac2002incremental, ozawa2008incremental, balsubramani2013fast] could be directly applied to OGD to compress the memory to a fixed size in an online manner. However, it typically involves repeated truncations, which can discard directions before their importance becomes apparent across tasks. Its choice of which principal components to discard, and thus, the quality of compression, is sensitive to the order in which the training data is encountered. The next section introduces an alternative means of approximation that avoids these issues.
II-C Matrix Approximation via Sketching
To reduce the memory cost of OGD, we use sketching, which approximates a matrix in lower dimensions. Our sketching methods update a fixed-size summary online as each new gradient arrives, independent of the total number of tasks or samples—crucial in continual learning scenarios with long, unknown horizons. They also enjoy linearity properties that make the final sketch equivalent to sketching the full gradients at once. Moreover, sketching is simple to implement and comes with well-developed approximation guarantees that we utilize in our analysis in Section IV.
@@ II-C Matrix Approximation via Sketching
Given a sketch $(\Omega,\Psi,Y,W)$ of a matrix $A\in\mathbb{R}^{m\times n}$, and letting $Q=\operatorname{orth}Y\in\mathbb{R}^{m\times\operatorname{rank}(Y)}$, $\widehat{A}\vcentcolon=Q(\Psi Q)^{\dagger}W$ approximates $A$, where $(\cdot)^{\dagger}$ denotes the pseudoinverse.
Rigorous justification for this is given by Tropp. The quality of the approximation improves with the size of the sketch set by $k$ and $l$. Especially, $k$ determines the rank of $\widehat{A}$.
@@
Given a sketch $(\Omega,\Psi,Y,W)$ of a symmetric matrix $A\in\mathbb{R}^{m\times m}$, let $Q=\operatorname{orth}Y\in\mathbb{R}^{m\times\operatorname{rank}(Y)}$ and define $X\vcentcolon=(\Psi Q)^{\dagger}W\in\mathbb{R}^{\operatorname{rank}(Y)\times m}$. Then, $\widehat{A}_{\mathrm{sym}}\vcentcolon=\tfrac{1}{2}(QX+X^{\top}Q^{\top})$ approximates $A$.
Formal justification is again given by Tropp. To approximate $\operatorname{range}A$, we can use the range of the concatenation $Q\;X^{\top}]$ from this proposition as we observe that:
@@
Algorithm 1 SketchOGD. The algorithm requires a model with weight vector $w$, a base optimizer such as SGD, as well as sketching subroutines from Sketching Method 1, 2 or 3.
@@
Sketching Method 1 Sketch of gradient matrix $G$.
@@
Sketching Method 2 Sketch of product $GG^{\top}$.
@@
Sketching Method 3 Symmetric sketch of product $GG^{\top}$.
@@
We note that the magnitudes of the gradients in $G$ play an important role in forming the approximations. In the case of full-memory OGD, i.e., when we have the luxury of storing all the directions of $G$, the magnitudes are redundant for preserving performance. However, under a fixed memory budget where storing all gradients is infeasible, the magnitudes provide crucial information for deciding which directions to retain. In particular, a larger magnitude of the gradient implies that a small model parameter update along that direction would alter the model output more significantly. Thus, the gradients with larger magnitudes become more important for mitigating overall catastrophic forgetting. Sketching methods naturally exploit this magnitude information by weighting directions accordingly.
III Sketched Orthogonal Gradient Descent
This section applies the matrix sketching methods introduced in Section II-C to obtain three variants of sketched orthogonal gradient descent (SketchOGD). We propose different variants as it is a priori unclear which performs best, but we will investigate this both by presenting theoretical bounds (Section IV-B) and by running experiments (Section VI). All variants adopt the following pattern:
  1. 1.
  2. 2.
  3. 3.
This procedure is described formally in Algorithm 1. Three variants are obtained by using either Sketching Method 1, 2 or 3 to supply the subroutines $\mathtt{InitializeSketch}$, $\mathtt{UpdateSketch}$ and $\mathtt{ExtractBasisFromSketch}$. The variants arise since there is freedom in both which matrix is sketched and what sketching technique is used. For instance, since $G$ and $GG^{\top}$ have the same range, either matrix may be sketched to approximate $\operatorname{range}G$. Furthermore, since $GG^{\top}$ is symmetric, Proposition II.3 may be applied. To elaborate, we propose three variants of SketchOGD:
SketchOGD-1: Algorithm 1 + Sketching Method 1. This algorithm directly sketches the gradient matrix $G$ via Proposition II.2, and $\operatorname{orth}Y$ for $Y=G\Omega$ forms the orthonormal basis. When a new gradient $g$ is received, we update $Y$ online instead of recomputing it from scratch since $Y^{(new)}=G^{(new)}\Omega^{(new)}=[G\;g]\begin{bmatrix}\Omega\\ \omega^{\top}\end{bmatrix}=Y+g\omega^{\top}$. Here, the random matrix $\Omega$ is incremented with a standard normal vector $\omega$ since the dimension of $G$ is changed. In terms of memory, SketchOGD-1 needs to store only $Y$ at a cost of $pk$.
SketchOGD-2: Algorithm 1 + Sketching Method 2. This algorithm sketches the product $GG^{\top}$ via Proposition II.2 as in SketchOGD-1. To update the sketch variable $Y=GG^{\top}\Omega$ online when a new gradient $g$ is received, we leverage the decomposition $Y^{(new)}=G^{(new)}{G^{(new)}}^{\top}\Omega=\sum_{c\in G}cc^{\top}\Omega=Y+gg^{\top}\Omega$ where the summation is running over the columns of $G$. In terms of memory, SketchOGD-2 needs to store both $Y$ and $\Omega$ at a cost of $2pk$.
SketchOGD-3: Algorithm 1 + Sketching Method 3. This algorithm sketches the product $GG^{\top}$ via symmetric sketch (Proposition II.3). The sketch variables $Y$ and $W$ are updated online using the decomposition of $GG^{\top}$ as in SketchOGD-2. The computation of the pseudoinverse $(\Psi Q)^{\dagger}$ is expedited by first taking the orthogonal-triangular ($\mathtt{QR}$) decomposition. The full sketch needs to be stored, at a memory cost of $2p(l+k)$. Since $l\geq k$, this is at least four times that of SketchOGD-1. However, the basis returned by Sketching Method 3 spans a higher-dimensional subspace.
@@ III Sketched Orthogonal Gradient Descent
As for computational complexities, SketchOGD-1,2 require $O(pk)$ for updating the sketch and $O(pk^{2})$ for extracting the orthonormal basis from the sketch via QR decomposition. For SketchOGD-3, updating the sketch costs $O(p(l+k))$. Extracting the basis involves QR decompositions of $Y$ and $\Psi Q$, and subsequent matrix multiplications to compute $X$. This yields a total cost on the order of $O(p(l+k)k)$. Note that $k$ and $l$ are chosen to be much smaller than the number of parameters $p$, so these additional costs are modest compared to the total training time dominated by backpropagation.
@@ IV-A Constructing a Metric Suited to SketchOGD
Given a matrix $G$ to be approximated and an orthonormal basis matrix $B$ returned by a sketching method, the reconstruction error is given by:
$\mathcal{E}_{G}(B)=\sum_{g\in G}\|g-\operatorname{proj}_{B}g\|_{2}^{2}=\|(I-BB^{\top})G\|_{F}^{2},$
where the summation is running over the columns of G, and $\operatorname{proj}_{M}$ for a matrix $M$ denotes the projection onto $\operatorname{range}M$. In SketchOGD, if the metric vanishes while $\operatorname{range}B\subseteq\operatorname{range}G$, then SketchOGD is equivalent to OGD.
@@
Another possible metric is the Grassmann distance between subspaces [grassmann], which could be applied to measure the distance between the subspace $\operatorname{range}G$ and the subspace $\operatorname{range}B$, where $B$ is returned by the sketching method. The drawback to this approach is that it neglects how the columns $\{g\in G\}$ are distributed within $\operatorname{range}G$. If certain directions appear more frequently, then it is more important for OGD that these directions are well-modeled.
@@ IV-B Deriving Bounds
We now provide a novel theoretical analysis of sketching applied to continual learning, through bounding our metric. While Tropp analyzed the Frobenius reconstruction error of sketching $\mathbb{E}\|A-\widehat{A}\|_{F}^{2}$, and Doan2020ATA bounded catastrophic forgetting, to our knowledge, we are the first to theoretically analyze the disruption a compression method causes to OGD.
@@
Define $U\in\mathbb{R}^{p\times p},\Sigma_{G}\in\mathbb{R}^{p\times n},V\in\mathbb{R}^{n\times n}$ as the singular value decomposition (SVD) of $G$ such that $G=U\Sigma_{G}V$, where $n$ is the number of sketched gradients. Then, we can define $\Sigma\vcentcolon=\Sigma_{G}\Sigma_{G}^{\top}\in\mathbb{R}^{p\times p}$ such that $GG^{\top}=U\Sigma U^{\top}$. Assuming singular values sorted in decreasing magnitude, we define $\Sigma_{1},\Sigma_{2},U_{1},U_{2}$ by splitting matrices at an index $\gamma\leq k-2$:
@@
Theoretical bounds on our metric based on this split would imply their dependence on the singular value decay of the gradients. In the basic sketching method where we directly sketch the matrix of gradients $G$, we can adapt a theorem of [halko] as below. See the appendix for the proof.
Theorem IV.1 (Expected error in Sketching Method 1).
Given that we create a sketch of $G$ with a parameter $k$ and obtain an orthonormal basis matrix $B$ in accordance with Sketching Method 1, and use the setup described in Definition IV.2, then the expected value of the metric $\mathcal{E}_{G}(B)$ has an upper bound:
$\mathbb{E}[\mathcal{E}_{G}(B)]\leq\min_{\gamma\leq k-2}(1+\tfrac{\gamma}{k-\gamma-1})\cdot\operatorname{Tr}{\Sigma_{2}}.$
@@
Note that sharper decays push the optimal split location $\gamma$ away from 0. For intuition, consider a flat spectral decay such that $\Sigma=\lambda I_{p\times p}$ for some $\lambda>0$ and the identity matrix $I_{p\times p}$. Then, $\gamma=0$ attains the minimum of the upper bound $p\lambda$. On the other hand, if $\Sigma=\Sigma_{G}\Sigma_{G}^{T}$ has linearly decaying spectral decay from $2\lambda$ to 0, then $\gamma=2(k{-}1){-}p$ attains the minimum approximately at $\frac{4\lambda}{p}(k{-}1)(p{-}k{+}1)$ for large enough $k$. Analysis can be found in the appendix.
We now provide two novel theorems for Sketching Method 2, one upper bounding the metric deterministically and the other in expectation. See the appendix for the proofs.
Theorem IV.2 (Deterministic error in Sketching Method 2).
We define $G,\Sigma_{1},\Sigma_{2},U_{1},U_{2}$ as in Definition IV.2 for a sketch parameter $k$ and split position $\gamma\leq k-2$. We create a sketch $(\Omega,\Psi,Y,W)$ of $A=GG^{\top}$ and obtain an orthonormal basis matrix $B$ in accordance with Sketching Method 2. We use $\Omega\in\mathbb{R}^{p\times k}$ to define: $\Omega_{1}:=U_{1}^{\top}\Omega$, $\Omega_{2}:=U_{2}^{\top}\Omega$. Assume $\Omega_{1}\in\mathbb{R}^{\gamma\times k}$ is full-rank. Then, if $\gamma$ is greater than the rank of $G$, then the metric $\mathcal{E}_{G}(B)=0$. Else,
$\mathcal{E}_{G}(B)\leq\|\Sigma_{2}\Omega_{2}\Omega_{1}^{\dagger}\Sigma_{1}^{-1/2}\|_{F}^{2}+\|\Sigma_{2}^{1/2}\|_{F}^{2}.$
@@
Theorem IV.3 (Exptected error in Sketching Method 2).
We define $G,\Sigma_{1},\Sigma_{2},U_{1}$ as in Definition IV.2 for a sketch parameter $k$ and arbitrary split position $\gamma\leq k-2$. We create a sketch $(\Omega,\Psi,Y,W)$ of $A=GG^{\top}$ and obtain an orthonormal basis matrix $B$ in accordance with Sketching Method 2. We assume $\Omega_{1}:=U_{1}^{\top}\Omega$ has full rank. Then, if $\gamma$ is greater than the rank of $G$, then the metric $\mathbb{E}[\mathcal{E}_{G}(B)]=0$. Else,
$\mathbb{E}[\mathcal{E}_{G}(B)]\leq\min_{\gamma\leq k-2}\tfrac{\gamma}{k-\gamma-1}\operatorname{Tr}(\Sigma_{2}^{2})\operatorname{Tr}(\Sigma_{1}^{-1})+\operatorname{Tr}(\Sigma_{2}).$